OpenStudy (anonymous):

PLEASE HELP! A plane took 5 hours to get from Omaha to San Francisco with a tailwind of 28 mph. It took 6 hours to make the return trip with a headwind of the same speed, 28 mph. How fast was the plane traveling in still air?

OpenStudy (anonymous):

1680miles

OpenStudy (anonymous):

how did you calculate that? are you sure?? Thanks!!!

OpenStudy (anonymous):

@satellite73 ?? do you know how to solve this??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and 1680 i think is too far

OpenStudy (anonymous):

i set up (5p+28)=6(p-28)..

OpenStudy (anonymous):

*5(p+28)

OpenStudy (anonymous):

close, i think it should be \[5(p+28)=6(p-28)\]

OpenStudy (anonymous):

if you call "p" the rate

OpenStudy (anonymous):

yeah what you said at first

OpenStudy (anonymous):

second

OpenStudy (anonymous):

Yeah but then when you simplify it's 28 and thats not an answer..

OpenStudy (anonymous):

oh i could be wrong, i jumped the gun maybe 1680 is the right answer

OpenStudy (anonymous):

but how do you get that?!?!?!

OpenStudy (anonymous):

lets solve \[5(x+28)=6(x-28)\] \[5x+140=6x-168\] \[140=x-168\] \[x=308\] so the rate is 308 miles per hour, but that is not the distance, that is the rate

OpenStudy (anonymous):

to find the distance replace the rate of 308 in either expression for example \[5(308+28)=5\times 336=1680\] as stated above

OpenStudy (anonymous):

THANK YOU!! but that is the rate in still air as well?

OpenStudy (anonymous):

@Anyssah was right , i jumped the gun

OpenStudy (anonymous):

the rate in still air is 308 mph, yes. the distance is therefore 1680

OpenStudy (anonymous):

Its not asking for the distance thought, its the rate it traveled in still air.

OpenStudy (anonymous):

ok, then 308 mph is the right answer to that question

OpenStudy (anonymous):

Thank you guys :)